Solve for velocity (in mph) and centripetal acceleration (in m/s^2) of a neutron star if its frequency is 500 Hertz and its radius is 12 kilometers
1
Expert's answer
2020-04-23T12:54:23-0400
"a=(2\pi f)^2r=(2\pi (500))^2(12000)=1.18\cdot10^{11}\frac{m}{s^2}"

"v=(2\pi f)r=(2\pi (500))(12000)=3.77\cdot10^{7}\frac{m}{s}\\=8.43\cdot10^{7}\ mph"